An increase in which of the following must increase the return on equity, all else constant? A. Net income and total equity B. Debt-equity ratio and total debt C. Total assets and sales D. Equity multiplier and total equity E. Total asset turnover and debt-equity ratio
Answer #1

ANS) E

Total Asset turnover and Debt Equity Ratio

it will increase Return on equity now lets see with the help of example

Case 1

sales 1000

asset 400

debt 100

equity 300

Asset Turn over = sales/ asset 1000/400 = 2.5

Debt Equity Ratio= 100/300= 0.33

And now if Debt and Sales will increase it will increase both Asset Turn over ration and Debt Equity ratio and increase in Debt will also give Leverage to company which will multiply its ROE(Return on equity)